# Class-Focused Auto-Reward Mechanism for Object Specific Text-to-Image models

### v1

@@ 1 Introduction
Self-rewarding is a concept that has been recently introduced in the field of Natural Language Processing (NLP) by W.Yuan et. al. Yuan et al. (2024) to address the limitations of traditional reward modeling based on reinforcement learning. More specifically, in Large Language Models (LLM), the reward function is typically defined by human annotators who rate the quality of generated text based on their preferences (Reinforcement learning from human feedback RLHF) Ouyang et al. (2022). This approach has several limitations, including the need for large amounts of human-annotated seed data, the potential for reward hacking (where the model learns to manipulate the reward function to generate higher-rewarding outputs), and the difficulty of accurately capturing human preferences in a reward function.
@@
Therefore, the self-reward pipeline proposed in Yuan et al. (2024) has the object of self-improving the LLM continuously through a model that can, not only generate text, but also self-judge its own content and improving the quality of the data. The authors suggest the LLM self-reward pipeline as follow: (1) start by generating a set of prompts using a pretrained model. (2) following this, N candidate responses are generated for each specific prompts by the same LLM. (3) the chosen responses is selected along with their prompts as ’preference pairs’. (4) the model is then trained using Direct Preference Optimization (DPO). DPO is an LLM training approach that has been proposed recently by Rafailov et. al. Rafailov et al. (2023). This method aims to align LLMs with human preferences by using a simple classification loss. DPO eliminates the need for sampling from the language model during the fine-tuning process. The main idea behind DPO is to optimize the likelihood of the preferred response over the dis-preferred response Amini et al. (2024). This is done by establishing the preferred and dis-preferred answers to the model for a specific prompt. The approach leads to an optimized likelihood of the preferred response and therefore resulting in a policy reformulation. DPO is simple, compatible with various language models architectures, and, unlike other LLM training approaches, does not require reinforcement learning or complex reward modeling.
@@
This paper explores the potential implementations of a concept analogous to self-improvement within the field of Text-to-Image (T2I) models. The overarching objective remains centered around similar rewarding concept to Yuan et al. (2024), wherein the T2I model not only refines its performance but actively engages in a self-improvement process. This transformative framework is characterized by a process of self-judgment and autonomous decision-making capabilities, introducing a novel approach to model evolution.
@@
Text-2-Image models, also known as diffusion models, are a class of generative models that learn to produce new images by simulating a diffusive process Zhang et al. (2023a). In the context of machine learning, this typically involves gradually adding noise to a data sample until the full image is covered with Gaussian Noise and then learning to reverse this process to generate new samples from the noise. This reverse generative process involves learning to transform the noisy samples into iteratively less and less noisy data. This can be performed using a neural network architecture that uses the information about diffused data and the current time step as inputs to output a prediction of the original data sample.
@@ 2 Method: Class-conditional Self-rewarding
The proposed self-rewarding, initially inspired from Yuan et al. (2024), works by combining the power of language performance and understanding of a LLM, text-to-image generation of image, and Image captioning also known as Image-to-text (I2T) to judge the generated images and provide the best possible dataset to T2I models. The self-rewarding pipeline could be summarized into 6 steps as shown in Figure 1.
@@ 2.1 T2I Prompts generation
Existing Language Model (LLM) architectures have demonstrated quiet impressive capabilities in understanding human language and even making contexts and emotions interaction Li et al. (2023), making them usable for various automation tasks Zhang et al. (2023b); Ozkaya (2023). Initially trained on massive textual datasets, those models have impressive text generation and completion capabilities and garnered attention for their proficiency in understanding nuanced language structures Wu et al. (2023). In the context of our research on Text-to-Image (T2I) diffusion models, leveraging the strengths of LLMs allows automating the generation of prompts.
To initiate the T2I Prompts generation process, the selection of an instruction-based fine-tuned LLM is required. Many LLM models have been fine-tuned for Q&A chat such as Llama 2 Touvron et al. (2023), Falcon series Almazrouei et al. (2023) and more recently Mistral Jiang et al. (2023) which seems to be efficient among low parameters models. Additionally, the advent of quantization allowed the efficient use of those models even on CPUs and low memory GPUs such as Activation-aware Weight Quantization (AWQ) Lin et al. (2023), Generative Pre-trained Transformers Quantization (GPTQ) Frantar et al. (2023). The selection of those quantized models is sufficient for the aforementioned process consisting of generating a list of prompts.
@@ 2.5 Training Text-to-Image model
LoRA Hu et al. (2021), short for Low-Rank Adaptation of Large Language Models, presents a pioneering approach devised by Microsoft researchers to efficiently fine-tune expansive language models. LoRA suggests a method wherein the pre-trained language model weights remain fixed, and trainable layers (referred to as rank-decomposition matrices) are introduced within each transformer block. This strategy significantly diminishes the count of trainable parameters and the GPU memory demands during training or inference, as gradients do not need to be computed for the majority of the model weights. By concentrating on the Transformer attention blocks of large-language models, the quality of fine-tuning achieved with LoRA matches that of full model fine-tuning, while being markedly swifter and less resource-intensive. When using LoRA, it is possible to avoid catastrophic forgetting of the text-to-image model abilities while outperforming plain fine-tuning Kumari et al. (2023).
LoRA differs from full-parameter fine-tuning in two fundamental parts: it tracks changes to weights instead of updating the weights entirely, and decomposes the weight matrix into two smaller matrices that contain the trainable parameters. Using decomposed matrices reduces the amount of GPU memory usage. By storing the model’s weights (called weights freezing Hu et al. (2021)) in memory during the training process, a lot of computation memory is saved. The model’s original weight matrix is left unchanged, only new weight matrices are summed to effectively fine-tune the language model. After training, the LoRA adapter matrices are then loaded to the retrained SD and later used for inference and evaluation. LoRA is accessed within the Hugging Face Parameter Efficient Fine-Tuning (PEFT) library Mangrulkar et al. (2022).HuggingFace library offers a convenient trainer for supervised finetuning with seamless integration for LoRA training and inference.
@@ 3.4.1 Self-judging Image-to-Text
To self-judge the quality of the images, we employed I2T for image captioning and analysis. The selected model for this task is Large Language And Vision Assistant (LlaVa) Liu et al. (2023) which have demonstrated it performance and have been proved to be efficient especially for describing and spotting unusual patterns in the images. To make the vision assistant LlaVa an image content-judging, the initially generated prompts from the LLM are given as an input along with the question about each individual image.
@@ 3.4.2 Open Vocabulary object detection filtering
In some cases, the generated images does not provide a clear view of the aimed class. As a filtering process, the selection of the optimal images out of the selected ones goes through an open vocabulary object detection model. For this process, we have selected the latest YOLO version called YOLO-world a real-time open vocabulary model Cheng et al. (2024).
@@ 3.5 Fine-tuning stable diffusion
When fine-tuning stable diffusion, it is necessary to preserve the visual features and general text-to-image capabilities of the base model on any prompt. However, in precise and retrained tasks, the goal of fine-tuning is to increase the precision and perceived accuracy of the model, while preserving the key stable diffusion abilities in text-to-image precision. For these reasons we use LoRA Hu et al. (2021) training as mentioned in 2.5. Using the filtered images and prompts pairs, the Stable Diffusion 2.1 Rombach et al. (2022) is retrained with LoRA in a traditional supervised fine-tuning fashion. The entire model retraining is done with the Diffusers library of HuggingFace von Platen et al. (2022). The training resolution for the images is set at 512, as this is the resolution of the original StableDiffusion 2.1, and the images are augmented with random horizontal flip. The training is completed on 2 NVIDIA A6000 48GB GPUs using mixed precision 16 float format, which allows memory optimization while ensuring no task-specific accuracy is lost compared to full precision training. The model is trained using LoRA for 100 epochs, with a batch size of 18, which is completed in two hours on the GPU hardware. The learning speed is set at a constant rate of 1e-04. Due to the nature of LoRA training and the selective dataset we are using, the inference results when applying the LoRA weights may vary. To evaluate the performances and changes, the additional weights are scaled between 0.0 and 1.0, translating to the intensity at which the LoRA weights are loaded. The most interesting changes occur between 0.2 and 0.8 LoRA weight scale, where the weights impact sufficiently the diffusion process without changing and losing the StableDiffusion creation entirely. Results of the inference at the LoRA scales 0.2, 0.4 and 0.7 are displayed in Figure 4. The relationship between prompts and generated images are evaluated using the CLIP score Radford et al. (2021), which provides a qualitative evaluation of the models, necessary for automation.
@@
Figure 4: Text-to-image compared results between the base model, our retrained StableDiffusion version at 0.4 and 0.7 LoRA weight intensity, and LimeWire Studio LimeWare for comparison. Below the generated images the CLIP score measures the compatibility of image-prompt pairs.
@@ 4 Discussion
To further evaluate the method’s efficiency in enhancing T2I models using validation prompts, the final model was assessed against the baseline SD Rombach et al. (2022) and SDXS Yuda Song (2024) (a fine-tuned version of SD). Each validation prompt was tested using a consistent PyTorch pseudo-random generation seed to ensure reproducibility. Our LoRA weights were adjusted across three different scales (0.2, 0.4, and 0.7) to examine their impacts on the SD checkpoints. The inference results were qualitatively evaluated using CLIP scoring Radford et al. (2021), similar to the LoRA training phase. The win-rate results for the 50 validation prompts, each tested with four different generation seeds, are illustrated in Fig. 3. In nearly all comparisons (with one exception pending results), our method demonstrated a higher win percentage against both the base and the SDXS models. Despite being close to 50%, the model frequently performed at least as well (wins + ties) as the baseline and fine-tuned SD, achieving an overall win-rate percentage of 70%.

### v2

# Class-Conditional self-reward mechanism for improved Text-to-Image models

@@ 1 Introduction
Self-rewarding is a concept that has been recently introduced in the field of Natural Language Processing (NLP) by [Yuan et al. (2024)] to address the limitations of traditional reward modeling based on reinforcement learning. More specifically, in Large Language Models (LLM), the reward function is typically defined by human annotators who rate the quality of generated text based on their preferences (Reinforcement learning from human feedback RLHF) [Ouyang et al. (2022)]. This approach has several limitations, including the need for large amounts of human-annotated seed data, the potential for reward hacking (where the model learns to manipulate the reward function to generate higher-rewarding outputs), and the difficulty of accurately capturing human preferences in a reward function.
@@
Therefore, the self-reward pipeline proposed in [Yuan et al. (2024)] has the object of self-improving the LLM continuously through a model that can, not only generate text, but also self-judge its own content and improving the quality of the data. The authors suggest the LLM self-reward pipeline as follow: (1) start by generating a set of prompts using a pretrained model. (2) following this, N candidate responses are generated for each specific prompts by the same LLM. (3) the chosen responses is selected along with their prompts as ’preference pairs’. (4) the model is then trained using Direct Preference Optimization (DPO). DPO is an LLM training approach that has been proposed recently by Rafailov et. al. [Rafailov et al. (2023)]. This method aims to align LLMs with human preferences by using a simple classification loss. DPO eliminates the need for sampling from the language model during the fine-tuning process. The main idea behind DPO is to optimize the likelihood of the preferred response over the dis-preferred response [Amini et al. (2024)]. This is done by establishing the preferred and dis-preferred answers to the model for a specific prompt. The approach leads to an optimized likelihood of the preferred response and therefore resulting in a policy reformulation. DPO is simple, compatible with various language models architectures, and, unlike other LLM training approaches, does not require reinforcement learning or complex reward modeling.
@@
This paper explores the potential implementations of a concept analogous to self-improvement within the field of Text-to-Image (T2I) models. The overarching objective remains centered around similar rewarding concept to [Yuan et al. (2024)], wherein the T2I model not only refines its performance but actively engages in a self-improvement process. This transformative framework is characterized by a process of self-judgment and autonomous decision-making capabilities, introducing a novel approach to model evolution.
@@
Text-2-Image models, also known as diffusion models, are a class of generative models that learn to produce new images by simulating a diffusive process [Zhang et al. (2023a)]. In the context of machine learning, this typically involves gradually adding noise to a data sample until the full image is covered with Gaussian Noise and then learning to reverse this process to generate new samples from the noise. This reverse generative process involves learning to transform the noisy samples into iteratively less and less noisy data. This can be performed using a neural network architecture that uses the information about diffused data and the current time step as inputs to output a prediction of the original data sample.
@@ 2 Method: Class-conditional Self-rewarding
The proposed self-rewarding, initially inspired from [Yuan et al. (2024)], works by combining the power of language performance and understanding of a LLM, text-to-image generation of image, and Image captioning also known as Image-to-text (I2T) to judge the generated images and provide the best possible dataset to T2I models. The self-rewarding pipeline could be summarized into 6 steps as shown in Figure 1.
@@ 2.1 T2I Prompts generation
Existing Language Model (LLM) architectures have demonstrated quiet impressive capabilities in understanding human language and even making contexts and emotions interaction [Li et al. (2023)], making them usable for various automation tasks [Zhang et al. (2023b); Ozkaya (2023)]. Initially trained on massive textual datasets, those models have impressive text generation and completion capabilities and garnered attention for their proficiency in understanding nuanced language structures [Wu et al. (2023)]. In the context of our research on Text-to-Image (T2I) diffusion models, leveraging the strengths of LLMs allows automating the generation of prompts.
To initiate the T2I Prompts generation process, the selection of an instruction-based fine-tuned LLM is required. Many LLM models have been fine-tuned for Q&A chat such as Llama 2 [Touvron et al. (2023)], Falcon series [Almazrouei et al. (2023)] and more recently Mistral [Jiang et al. (2023)] which seems to be efficient among low parameters models. Additionally, the advent of quantization allowed the efficient use of those models even on CPUs and low memory GPUs such as Activation-aware Weight Quantization (AWQ) [Lin et al. (2023)], Generative Pre-trained Transformers Quantization (GPTQ) [Frantar et al. (2023)]. The selection of those quantized models is sufficient for the aforementioned process consisting of generating a list of prompts.
@@ 2.5 Training Text-to-Image model
LoRA [Hu et al. (2021)], short for Low-Rank Adaptation of Large Language Models, presents a pioneering approach devised by Microsoft researchers to efficiently fine-tune expansive language models. LoRA suggests a method wherein the pre-trained language model weights remain fixed, and trainable layers (referred to as rank-decomposition matrices) are introduced within each transformer block. This strategy significantly diminishes the count of trainable parameters and the GPU memory demands during training or inference, as gradients do not need to be computed for the majority of the model weights. By concentrating on the Transformer attention blocks of large-language models, the quality of fine-tuning achieved with LoRA matches that of full model fine-tuning, while being markedly swifter and less resource-intensive. When using LoRA, it is possible to avoid catastrophic forgetting of the text-to-image model abilities while outperforming plain fine-tuning Kumari et al. (2023).
LoRA differs from full-parameter fine-tuning in two fundamental parts: it tracks changes to weights instead of updating the weights entirely, and decomposes the weight matrix into two smaller matrices that contain the trainable parameters. Using decomposed matrices reduces the amount of GPU memory usage. By storing the model’s weights (called weights freezing [Hu et al. (2021)]) in memory during the training process, a lot of computation memory is saved. The model’s original weight matrix is left unchanged, only new weight matrices are summed to effectively fine-tune the language model. After training, the LoRA adapter matrices are then loaded to the retrained SD and later used for inference and evaluation. LoRA is accessed within the Hugging Face Parameter Efficient Fine-Tuning (PEFT) library [Mangrulkar et al. (2022)].HuggingFace library offers a convenient trainer for supervised finetuning with seamless integration for LoRA training and inference.
@@ 3.4.1 Self-judging Image-to-Text
To self-judge the quality of the images, we employed I2T for image captioning and analysis. The selected model for this task is Large Language And Vision Assistant (LlaVa) [Liu et al. (2023)] which have demonstrated it performance and have been proved to be efficient especially for describing and spotting unusual patterns in the images. To make the vision assistant LlaVa an image content-judging, the initially generated prompts from the LLM are given as an input along with the question about each individual image.
@@ 3.4.2 Open Vocabulary object detection filtering
In some cases, the generated images does not provide a clear view of the aimed class. As a filtering process, the selection of the optimal images out of the selected ones goes through an open vocabulary object detection model. For this process, we have selected the latest YOLO version called YOLO-world a real-time open vocabulary model [Cheng et al. (2024)].
@@ 3.5 Fine-tuning stable diffusion
When fine-tuning stable diffusion, it is necessary to preserve the visual features and general text-to-image capabilities of the base model on any prompt. However, in precise and retrained tasks, the goal of fine-tuning is to increase the precision and perceived accuracy of the model, while preserving the key stable diffusion abilities in text-to-image precision. For these reasons we use LoRA [Hu et al. (2021)] training as mentioned in 2.5. Using the filtered images and prompts pairs, the Stable Diffusion 2.1 [Rombach et al. (2022)] is retrained with LoRA in a traditional supervised fine-tuning fashion. The entire model retraining is done with the Diffusers library of HuggingFace [von Platen et al. (2022)]. The training resolution for the images is set at 512, as this is the resolution of the original StableDiffusion 2.1, and the images are augmented with random horizontal flip. The training is completed on 2 NVIDIA A6000 48GB GPUs using mixed precision 16 float format, which allows memory optimization while ensuring no task-specific accuracy is lost compared to full precision training. The model is trained using LoRA for 100 epochs, with a batch size of 18, which is completed in two hours on the GPU hardware. The learning speed is set at a constant rate of 1e-04. Due to the nature of LoRA training and the selective dataset we are using, the inference results when applying the LoRA weights may vary. To evaluate the performances and changes, the additional weights are scaled between 0.0 and 1.0, translating to the intensity at which the LoRA weights are loaded. The most interesting changes occur between 0.2 and 0.8 LoRA weight scale, where the weights impact sufficiently the diffusion process without changing and losing the StableDiffusion creation entirely. Results of the inference at the LoRA scales 0.2, 0.4 and 0.7 are displayed in Figure 4. The relationship between prompts and generated images are evaluated using the CLIP score [Radford et al. (2021)], which provides a qualitative evaluation of the models, necessary for automation.
@@
Figure 4: Text-to-image compared results between the base model, our retrained StableDiffusion version at 0.4 and 0.7 LoRA weight intensity, and LimeWire Studio [LimeWare ] for comparison. Below the generated images the CLIP score measures the compatibility of image-prompt pairs.
@@ 4 Discussion
To further evaluate the method’s efficiency in enhancing T2I models using validation prompts, the final model was assessed against the baseline SD [Rombach et al. (2022)] and SDXS [Yuda Song (2024)] (a fine-tuned version of SD). Each validation prompt was tested using a consistent PyTorch pseudo-random generation seed to ensure reproducibility. Our LoRA weights were adjusted across three different scales (0.2, 0.4, and 0.7) to examine their impacts on the SD checkpoints. The inference results were qualitatively evaluated using CLIP scoring [Radford et al. (2021)], similar to the LoRA training phase. The win-rate results for the 50 validation prompts, each tested with four different generation seeds, are illustrated in Fig. 3. In nearly all comparisons (with one exception pending results), our method demonstrated a higher win percentage against both the base and the SDXS models. Despite being close to 50%, the model frequently performed at least as well (wins + ties) as the baseline and fine-tuned SD, achieving an overall win-rate percentage of 70%.